The sound level produced by one singer is
84.2 dB.
What would be the sound level produced
by a chorus of 43 such singers (all singing at
the same intensity at approximately the same
distance as the original singer)?
Answer in units of dB.
1
Expert's answer
2020-02-18T10:16:08-0500

If the intensity of sound produced by one singer is I, then the intensity of sound of 43 singers is 43I. In decibels, this intensity will be equal to:


"X=10\log_{10}\frac{43I}{I_0}"

"X=10\log_{10}40 +10\log_{10}\frac{I}{I_0}=10\log_{10}40 +X_0"

"X=10\log_{10}40 +84.2=100.2\ dB"